# WIR3D: Visually-Informed and Geometry-Aware 3D Shape Abstraction

### v1

@@ 1 Introduction
To achieve our sparse and visually salient abstraction, we leverage spatially-conditioned semantic supervision by utilizing the intermediate activations of a 2D pre-trained foundation model (CLIP [47]). Our choice of model is justified by prior work demonstrating CLIP’s strong performance in tasks requiring high-level abstract understanding across different modalities [16, 60, 18]. We observe that both different layers and architectures enable different levels of control over geometric and texture elements of the shape and thus split our optimization into separate geometry and texture abstraction stages.
Our spatial conditioning comes from 3D keypoints, which determine the weight that specific visual features are given in our abstraction. The keypoints being in 3D ensures their multi-view consistency and also enables fine-grained user control over the resulting abstraction. We show in Fig. 9 an application where users can iteratively add detail to the abstraction through selecting keypoints corresponding to salient features.
We further encourage curve adherence to the input geometry using a neural SDF loss. The surface compliance enables an application where the curves can be used as intuitive deformation handles for the shape. We demonstrate this application in Fig. 8 and the supplemental, where our curves’ effectiveness as control handles stems from their alignment to semantically meaningful regions of the surface and texture.
@@ 3.1 Curve Representation
The resulting image from this projection and rasterization process is $I_{\text{curve}}=\mathcal{R}(\{B_{i}\})$. We refer to target shape renders as $I_{\text{target}}$.
@@ 3.2 Losses
$+\sum_{l=3,4}||\text{CLIP}_{l}(I_{\text{curve}})-\text{CLIP}_{l}(I_{\text{target}})||_{2}^{2}$
@@
$I_{\text{weight}}(x,y)=1+\sum_{p}e^{\frac{-||(x,y)-p||^{2}}{2\sigma^{2}}}$ (3)
where $x,y\in[0,1]$ index the normalized image pixels, and $p\in[0,1]$ is the keypoint projected to normalized coordinates. We add 1 to the weights so that regions far from the keypoints still contribute to the loss. We use these constructed weight maps to weight our semantic loss Eq. 2 for each render, such that each L2 term in the intermediate layer losses $||\text{CLIP}_{l}(I_{\text{curve}})-\text{CLIP}_{l}(I_{\text{target}})||_{2}^{2}$ becomes $||I_{\text{weight}}\cdot(\text{CLIP}_{l}(I_{\text{curve}})-\text{CLIP}_{l}(I_{\text{target}}))||_{2}^{2}$. The weight map is downsampled to match the resolution of each intermediate feature map. We visualize this process in Fig. 4.
@@
$+\sum_{l=3,4}||I_{\text{weight}}\cdot(\text{CLIP}_{l}(I_{\text{curve}})-\text{CLIP}_{l}(I_{\text{target}}))||_{2}^{2}$
@@
A natural question is how sensitive this loss is to poor keypoint selection. We show an example optimization using our localization loss with randomly distributed keypoints in Fig. 18 the supplemental, and show the result is not much different than optimizing without any keypoints ($\mathcal{L}_{\text{semantic}}$). We also ablate on $\mathcal{L}_{\text{local}}$ when keypoints identify salient features in Fig. 17, and show the localization weighting is critical for abstracting fine-grained features. This demonstrates that $\mathcal{L}_{\text{local}}$ is a strict improvement over $\mathcal{L}_{\text{semantic}}$ in cases where the keypoints are meaningful, and otherwise produces robust results on par with $\mathcal{L}_{\text{semantic}}$.
On top of the localized keypoint loss, we include SDF and view regularization losses to ensure the stroke abstraction is represented in all possible views.
@@
We further regularize the abstraction by enforcing that all curves are visible from all sampled viewing angles. This forces all curves to participate in the shape abstraction from every angle, which is necessary for a proper 3D abstraction. With $\mathcal{P}$ indicating the perspective projection of the Bézier curve control keypoints, we have:
$\mathcal{L}_{\text{ndc}}$ $=\sum_{i=1}^{n}\sum_{j=1}^{2}\text{ReLU}(\mathcal{P}(B_{i})_{j}-1)+\text{ReLU}(-\mathcal{P}(B_{i})_{j}),$ (6)
where ReLU is a Rectified Linear Unit.
@@ 3.3 Two-Stage Optimization
In our experiments we find that different CLIP architectures and layer combinations are sensitive to geometry or semantic shape features. To exploit this, we construct a two-stage training pipeline, in which the first stage is optimizes for the shape geometry and the second stage abstracts the shape texture, where each stage leverages different CLIP architectures and semantic losses.
@@
$\mathcal{L}_{\text{II }}=$ $\mathcal{L}_{\text{local}}(I_{\text{curve}},I_{\text{target}}^{\text{surface}})+\mathcal{L}_{\text{SDF}}+\mathcal{L}_{\text{ndc}}(\{B_{i}\}).$
We use CLIP architectures RN101 and RN50x16 for $\mathcal{L}_{\text{semantic}}$ and $\mathcal{L}_{\text{local}}$, respectively. We’ve found empirically that RN101 tends to be more sensitive to geometric structures, whereas RN50x16 is more sensitive to higher-level visual concepts.
@@ 4 Experiments
For all experiments, we randomly sample views spanning 0 to 30 degrees elevation and 0 to 360 degrees azimuth.
@@ 4.1 Qualitative Results
Fig. 5 shows 3D stroke abstractions generated by WIR3D for textured shapes.WIR3D also successfully abstracts textures and captures structured patterns, such as watermelon seeds (Row 1) and spots (Rows 5/6 Spot/Bob), as well as facial features (Rows 5/7 Spot/Nefertiti). We also show results on untextured shapes in Fig. 13 in the supplemental. Our method can represent complex geometries, such as the spiral column band (Row 6), the parallel rows of spines on the stegosaurus (Row 5).
@@
Multi-View Fidelity.
Because our curves are defined in 3D, our abstraction is view-consistent by construction. However, this does not guarantee the curves plausibly represent the shape from arbitrary views. Fig. 14 in the supplemental shows that our abstraction faithfully represents the shape for densely sampled views in a 360 range.
@@ 4.2 Baseline Comparisons
The baselines against which we compare are NEF [67] and 3Doodle [7]. For both methods, we use the publicly available repositories published by the respective authors. We standardize the sampled views during optimization, and use the default hyperparameters otherwise. 3Doodle originally initializes its curves using SFM, but given the method’s sensitivity to poor initialization, we standardize initialization to the same furthest point samples as our method, and find that they perform better than SFM.
@@
We show qualitative comparisons with the 3Doodle baseline for both untextured and textured shapes in Fig. 6. Our method consistently captures the global geometry, whereas 3Doodle struggles in a low signal context such as untextured shapes (missing back left chair leg, flattened bird geometry). Furthermore, 3Doodle is consistently insensitive to facial structures, whereas our localized weighting ensures that we abstract them (rows 2, 6). Similarly, we are able to represent fine-grained texture structures as high-level visual motifs, such as the dragon scales (row 3).
Figure 6: Qualitative comparison. We compare a subset of our textured and untextured results against 3Doodle. 3Doodle produces reasonable abstractions that capture the overall structure but lack precision when it comes to the feature details and specific geometric structures.
Quantitative Comparison.
We assess the quality of the stroke abstractions using three perceptual metrics and one geometric metric, reported in Tab. 1. The first two, LPIPS [68] and $\text{CLIP}^{\text{img}}$ [61], are the same metrics reported by 3Doodle. Our method outperforms 3Doodle on both metrics, though numerically our $\text{CLIP}^{\text{img}}$ is not much higher. We note that CLIP has been shown to be insensitive to large visual differences, as systematized in [56, 63]. We further illustrate this point in Fig. 12 (supplemental), where we show that a randomly oriented cow outline obtains a similar $\text{CLIP}^{\text{img}}$ score as the edge map [3] of the ground-truth render. Thus, the numerical difference in $\text{CLIP}^{\text{img}}$ is an unreliable indicator of the difference in visual quality.
This motivates our user perceptual study ($N=96$), which takes all the results from our dataset and compares our Wir3D optimized strokes against the 3Doodle optimized strokes. Specifically, we display rotating gifs of both results side by side, along with a rotating gif of the target shape, and ask users to rank the 3D strokes based on how well they represent the target shape. We collected responses from 96 users and compute the frequency our method is ranked higher than 3Doodle. Our curves are ranked as better shape abstractions 88% of the time. Screenshots from the study are shown in the supplemental.
Our final metric, “Coverage”, is a geometric metric that quantifies how well the stroke abstractions cover the original surface. We aim to abstract an input 3D model, so fidelity to the input geometry is key to abstraction quality. We measure coverage by sampling 100k points over the input surface and computing the 1-direction Chamfer distance from the surface points to the optimized strokes. This metric evaluates whether each point on the surface has a curve reasonably close to it. Adequate coverage of a shape’s geometry is important for downstream applications which make use of the surface correspondence, such as the deformation application shown in Fig. 8. Our method significantly outperforms the baselines in terms of coverage, with a $>$2x reduction in coverage distance relative to 3Doodle, and $>$4x reduction relative to NEF.
@@
Table 1: We compare between WIR3D, 3Doodle [7], and NEF [67] through 3 perceptual metrics computed over novel views of the curves and the target shape. “LPIPS” measures the average LPIPS similarity score using AlexNet. “$\text{CLIP}^{\text{img}}$” measures the average cosine similarity scaled between (0,1) using the ViT/B-32 model. “User Rank” measures the percentage of user responses that prefer one method over the other ($N=96$). We also compute a geometric metric “Coverage”, which measures the 1-direction Chamfer distance between the curves and the target 3D surface.
@@ 4.3 Ablations
We perform a thorough ablation study demonstrating the importance of our design choices in WIR3D. Figures for each ablation are shown in the supplemental. We report the ablation metrics in Tab. 2. Removing the CLIP intermediate layers from our supervision results in the highest reduction in quality, followed by removing the stage 1 training. Removing the localized keypoint loss also results in some reduction in the metrics, but as discussed in Sec. 4.2, perceptual metrics may not be sensitive to the fine visual detail this loss is designed to capture. Removing the SDF loss results in worse geometric coverage, as expected.
Stage 1.
Stage 1 training ensures the full geometry is represented from every view. Removing this stage results in the optimization biasing towards certain views and creating an overall “flattened” effect of the geometry (Fig. 15).
No intermediate CLIP layers.
Localized keypoint loss.
As discussed in Sec. 3.2, our method without keypoints produces reasonable abstraction but lacks fine-grained detail (Fig. 17). We also show in Fig. 18 that initialization with random keypoints is no worse than running our method without keypoints. Hence, we are robust to poor keypoint choice.
SDF loss.
Our SDF loss prevents semantic features from pulling off the surface implied by the geometry curves. It also prevents occasional small floaters (Fig. 19).
@@ 4.4 Applications
We demonstrate two applications enabled by our 3D curve representation. The first is user-interactive feature control. Our localized weighting framework allows for user control over which features are represented in the abstraction. After optimization, the user can further refine the curves by selecting additional keypoints, and new curves can be quickly optimized to add detail to the feature of interest. We demonstrate this refinement procedure in Fig. 9, where keypoints can be used to make specific structures more explicit (e.g. wheels on the plane) or to add texture detail (e.g. nefertiti headband). The refinement is rapid, and is completed in a few hundred iterations, or around a minute.
The curves’ adherence to the input geometry makes them intuitive deformation handles. We show in Fig. 8 examples of interactive deformation using the curves, where Euclidean distance-based skinning weights are used to map deformations from the curves to the shape surface. The curves naturally wrap salient features so that the deformations are smoothly transferred, and relevant parts are easily manipulated. More detail and videos are in the supplemental.
@@ 5 Conclusion
Limitations. The quality of WIR3D is primarily limited by the quality of the detected/input keypoints. In cases where the keypoints are located in non-semantic regions, our abstraction will perform on par with the result without localized weighting. Another limitation is the preprocessing required for our method. Fitting a neural SDF, running our automatic keypoint algorithm, and generating the freestyle renders may take a lot of time depending on the complexity of the input model. For instance, the total preprocessing time for our Nefertiti model (100,000 faces) is around 2 hours. Future work could look into making this preprocessing more efficient or removing it entirely, and producing better algorithms for semantic keypoint detection.
@@ A Data Preprocessing
Figure 11: Freestyle render ablation. Running our method without freestyle renders still produces a reasonable abstraction, but key geometric features, such as the wheels of the car, may be missed due to the lack of visual content from the surface renders.
@@ B Neural Edge Field Comparison
Figure 13: Qualitative results for untextured shapes. We show the result of our method on a collection of untextured meshes. Our method is effective and robust on a wide collection of different geometries.
@@
Figure 17: Localized keypoint ablation. We show the influence of our localized weighting framework when keypoints pinpoint semantically meaningful features. The result without keypoints is reasonable, but the face texture details are only captured with the introduction of keypoints for spatial weighting.
Figure 18: Random keypoints ablation. We compare our method with randomly sampled keypoints to our method with no keypoints (unweighted semantic loss). Note that the result with random keypoints is of similar quality to our method without keypoints, which demonstrates that our method is robust to non-meaningful keypoints.
@@ D Ablations
Random Keypoints Ablation.
We ablate on the keypoint selection by running our method with the localized keypoint loss weighted by random keypoints sampled from the surface and compare the results to our method run without keypoint weighting (unweighted semantic loss). The results are in Fig. 18. Note that the quality of our method with random keypoints is the same as running our method with no keypoints, demonstrating our robustness to non-meaningful keypoints. Our localized keypoint loss is strictly better than the unweighted semantic loss when the keypoints are meaningful, and on par otherwise.

### v2

@@ 1 Introduction
To achieve our abstraction, we leverage spatially-conditioned semantic supervision by utilizing the intermediate activations of a 2D pre-trained foundation model (CLIP [48]). Our choice of model is justified by previous work that demonstrates CLIP’s strong performance in tasks requiring high-level abstract understanding across different modalities [16, 62, 18]. We observe that both different layers and architectures enable different levels of control over geometry and texture elements, so we split our optimization into separate geometry and texture abstraction stages.
Our spatial conditioning comes from 3D keypoints, which determine the weight specific visual features are given in our abstraction. The keypoints being in 3D ensures their multi-view consistency and enables fine-grained user control. We show in Fig. 11 an application where users can iteratively add detail to the abstraction through selecting keypoints corresponding to salient features.
We encourage curve adherence to the input geometry using a neural SDF loss. The surface compliance enables an application where the curves are used as intuitive deformation handles, shown in Fig. 10 and the supplemental. Our curves’ effectiveness as control handles stems from their alignment to semantically meaningful regions of the surface and texture.
@@ 3.1 Curve Representation
The resulting image from this rasterization process is $I_{\text{curve}}=\mathcal{R}(\{B_{i}\})$. We refer to target shape renders as $I_{\text{target}}$.
@@ 3.2 Losses
[…]
@@
[…]
where $x,y\in[0,1]$ index the normalized image pixels, and $p\in[0,1]$ is the keypoint projected to normalized coordinates. We add 1 to the weights so that regions far from the keypoints still contribute to the loss. We use these constructed weight maps to weight our semantic loss Eq. 2 for each render, such that each L2 term in the intermediate layer losses $\text{CLIP}_{l}(I_{\text{curve}})-\text{CLIP}_{l}(I_{\text{target}})||_{2}^{2}$ becomes $I_{\text{weight}}\cdot(\text{CLIP}_{l}(I_{\text{curve}})-\text{CLIP}_{l}(I_{\text{target}}))||_{2}^{2}$. The weight map is downsampled to match the resolution of each intermediate feature map. We visualize this process in Fig. 4.
@@
[…]
@@
We further regularize the abstraction by enforcing that all curves are visible from all sampled viewing angles, which prevents curves from hiding outside of the camera view. With $\mathcal{P}$ indicating the perspective projection of the Bézier curve control keypoints, we have:
$\mathcal{L}_{\text{ndc}}$ $=\sum_{i=1}^{n}\sum_{t}\text{ReLU}(\mathcal{P}(B_{i}(t))-1)+\text{ReLU}(-\mathcal{P}(B_{i}(t)))$ (6)
where ReLU is a Rectified Linear Unit. This penalizes all curve points with image coordinates outside of the render frame (0-1). The first term has non-zero value when the coordinates are greater than 1, and the second term has non-zero value when the coordinates are less than 0.
@@ 3.3 Two-Stage Optimization
In our experiments we find that different CLIP architectures are sensitive to geometry or semantic shape features. To exploit this, we use a two-stage training pipeline, in which the first stage optimizes for the shape geometry and the second stage abstracts the shape texture, and each stage leverages different CLIP architectures and semantic losses.
@@
[…]
We use CLIP architectures RN101 and RN50x64 for $\mathcal{L}_{\text{semantic}}$ and $\mathcal{L}_{\text{local}}$, respectively. We’ve found empirically that RN101 tends to be more sensitive to geometric structures, whereas RN50x64 is more sensitive to higher-level visual concepts.
@@ 4 Experiments
For all experiments, we randomly sample views spanning 0 to 30 degrees elevation and 0 to 360 degrees azimuth. We use 20 curves for stage 1 and initialize 20 additional curves for stage 2.
@@ 4.1 Qualitative Results
Fig. 5 shows 3D stroke abstractions generated by WIR3D for textured shapes.WIR3D successfully abstracts textures and small salient structures, such as watermelon seeds (Row 1) and spots (Rows 5/6), as well as facial features (Rows 5/7). We show results on untextured shapes in Fig. 15 in the supplemental, which demonstrate that our method can also represent complex geometries.
@@
In-The-Wild Abstraction.
WIR3D has no requirements on the quality of the input geometry. Only the Freestyle rendering will be affected by poor geometry, but CLIP supervision is robust to noisy render artifacts due to its global semantic understanding. We show in Fig. 7 that WIR3D successfully abstracts shapes built from photo reconstruction software, which have many geometric defects.
Figure 7: In-The-Wild Reconstructions. We abstract two shapes from the Real-World-Textured Things dataset [36], which provides textured models reconstructed using modern off-the-shelf photo-reconstruction tools on in-the-wild images. The shoe (top) contains 360 boundary edges, 31 non-manifold edges, and 9 non-manifold vertices. The sculpture (bottom) contains 113601 boundary edges and 1274 non-manifold edges.
@@ 4.2 Baseline Comparisons
The baselines we compare to are NEF [69] and 3Doodle [7]. For both methods, we use the publicly available repositories published by the respective authors. We standardize the sampled views and use the default hyperparameters otherwise. 3Doodle initializes its curves using SFM, but given the method’s sensitivity to poor initialization, we standardize initialization to the same furthest point samples as our method, finding that they perform better than SFM. We use 40 curves for 3Doodle to match WIR3D, but NEF has no ability to determine the curve count.
@@
Fig. 8 compares against WIR3D against 3Doodle for both textured and untextured shapes. Our method consistently captures the global geometry, whereas 3Doodle struggles in a low signal context such as untextured shapes (missing back left chair leg, flattened bird geometry). 3Doodle is consistently insensitive to facial structures, whereas our localized weighting ensures that we abstract them (rows 2, 6). Similarly, we are able to represent fine-grained texture structures as high-level visual motifs, such as the dragon scales (row 3).
Figure 8: Qualitative comparison. We compare a subset of our textured and untextured results against 3Doodle. 3Doodle produces reasonable abstractions that capture the overall structure but lack precision when it comes to the feature details.
@@
Table 1: We compare WIR3D, 3Doodle [7], and NEF [69] through 3 perceptual metrics computed over novel views of the curves and the target shape. “LPIPS” measures the average LPIPS similarity score using AlexNet. $\text{CLIP}^{\text{img}}$” measures the average cosine similarity scaled between (0,1) using the ViT/B-32 model. “User Rank” measures the percentage of user responses that prefer one method over the other ($N=96$). We also compute a geometric metric “Coverage”, which measures the 1-direction Chamfer distance between the curves and the target 3D surface.
Quantitative Comparison.
We use three perceptual metrics and one geometric metric, reported in Tab. 1. The first two, LPIPS [70] and $\text{CLIP}^{\text{img}}$ [63], are the same metrics reported by 3Doodle. Our method outperforms 3Doodle on both metrics, though numerically our $\text{CLIP}^{\text{img}}$ is not much higher. We note that CLIP has been shown to be insensitive to large visual differences, as systematized in [58, 65]. We further illustrate this point in Fig. 14 (supplemental), where we show that a randomly oriented cow outline obtains a similar $\text{CLIP}^{\text{img}}$ score as the edge map [3] of the ground-truth render. Thus, the numerical difference in $\text{CLIP}^{\text{img}}$ is an unreliable indicator of the difference in visual quality.
This motivates our user perceptual study ($N=96$), which takes all the results from our dataset and compares Wir3D results against 3Doodle. Specifically, we display rotating gifs of both results side by side, along with a rotating gif of the target shape, and ask users to rank the 3D strokes based on how well they represent the target shape. We collected responses from 96 users and compute the frequency our method is ranked higher than 3Doodle. Our curves are ranked higher 88% of the time. Screenshots from the study are shown in the supplemental.
Our final metric, “Coverage”, is a geometric metric that quantifies how well the stroke abstractions cover the original surface. We measure coverage by sampling 100k surface points and computing the 1-direction Chamfer distance from the points to the strokes. This evaluates whether each point on the surface has a curve reasonably close to it. Our method significantly outperforms the baselines in terms of coverage, with a >>2x reduction in coverage distance relative to 3Doodle, and >>4x reduction relative to NEF.
@@ 4.3 Ablations
We perform a thorough ablation study demonstrating the importance of our design choices in WIR3D. Figures for each ablation are shown in the supplemental. We report the ablation metrics in Tab. 2. Removing the CLIP intermediate layers (“No CLIP Layers”) from our supervision results in the highest reduction in quality (supplemental Fig. 18), followed by removing the stage 1 training (“w/o Stage 1”). Removing this stage results in the optimization biasing towards certain views and creating an overall “flattened” effect of the geometry (supplemental Fig. 17). Removing the localized keypoint loss (“No Local”) also results in some reduction in the metrics, but as discussed in Sec. 4.2, perceptual metrics may not be sensitive to the fine visual detail this loss is designed to capture. Removing the SDF loss results in worse geometric coverage, as expected (“No SDF”).
Keypoint Robustness.
We test the robustness of our keypoint loss to noisy keypoints. In Fig. 9, we randomly sample keypoints and compare against the original results (“With Keypoints”) and the results without using the localization weighting (“No Keypoints”). When noisy keypoints are given, the results are similar to the results without localization weighting (becomes Eq. 2). With meaningful keypoints, our method produces strictly improved results.
Figure 9: Keypoint Robustness. Our localized keypoint loss is robust. When keypoints are random (“Random Keypoints”; positions in insets shown), the result is similar to when the localized keypoint loss is not used at all (“No Keypoints”).
@@ 4.4 Applications
We demonstrate two applications enabled by our 3D curve representation. The first is user-interactive feature control. By selecting additional keypoints, new curves can be added and quickly optimized to add detail to any feature of interest. We show in Fig. 11 that keypoints can be used to make structures more explicit (e.g. wheels on the plane) or to add texture detail (e.g. nefertiti headband). The refinement is completed in a few hundred iterations (1 min).
The curves’ adherence to the input geometry makes them intuitive deformation handles. Fig. 10 shows examples of interactive deformations, where Euclidean distance-based skinning weights map deformations from the curves to the shape surface. The curves naturally wrap salient features so that the deformations are smoothly transferred, and relevant parts are easily manipulated. To verify the effectiveness of these curves as deformation handles, we conduct a perceptual study comparing deformations generated by WIR3D against ARAP [53], a classic deformation method. We drag a single control point on a WIR3D curve and compare the result to dragging the nearest vertex with ARAP. We ask respondents to “pretend to be 3D modelers in charge of sculpting various shapes”, and ask them which deformation result is most desirable from that perspective.
The respondents (N=42) found our WIR3D-based deformations to be more desirable 80% of the time. We show one example in the inset, with remaining study details in the supplemental. Our curve-based handles preserve the underlying shape semantics better, such as deforming only the cow’s horn while leaving its head intact.
@@ 5 Conclusion
Limitations. The quality of WIR3D is conditional on how good the keypoints are (Fig. 9). Another limitation is the preprocessing. Fitting a neural SDF, running our automatic keypoint algorithm, and generating the Freestyle renders may take a lot of time depending on the input (2 hours for Nefertiti at 100k faces).
6 Acknowledgements
This research was supported by grant #2022363 from the United States - Israel Binational Science Foundation (BSF), grant #2304481 from the National Science Foundation (NSF), and gifts from Adobe, Snap, and Google.
@@ A Data Preprocessing
Figure 13: Freestyle render ablation. Running our method without freestyle renders still produces a reasonable abstraction, but key geometric features, such as the wheels of the car, may be missed due to the lack of visual signal from the surface renders.
@@ B Neural Edge Field Comparison
Figure 15: Qualitative results for untextured shapes. We show the result of our method on a collection of untextured meshes. Our method is effective and robust on a wide collection of different geometrie, such as the spiral column band (Row 6), the parallel rows of spines on the stegosaurus (Row 5).
@@ D Ablations
No intermediate CLIP layers.
No Stage 1.
Stage 1 is essential for obtaining abstractions with visual volume. Without it the abstractions are flattened, so look reasonable from certain angles but not in others (Fig. 17).